OpenStudy (anonymous):

What is the simplified form of 1 over x–cubed all over 3 over x squared ?

OpenStudy (anonymous):



OpenStudy (anonymous):

3x?

OpenStudy (anonymous):

\[\frac{\frac {a}{b}}{\frac {c}{d}} = \frac{a \times d}{b \times c}\]

jhonyy9 (jhonyy9):

1/x^3 * x^2/3 = 1/3x

OpenStudy (anonymous):

ok but wouldnt the one just go away?

OpenStudy (anonymous):

like u would know its there regardless

OpenStudy (anonymous):

False. Only if the one is on the bottom will it go away.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Aka (3x)/1 = 3x, but 1/(3x) = 1/(3x) and that's all it can be.

jhonyy9 (jhonyy9):

so if you need divide two fractions you need multiply first by second inversed - this is by definition